OpenStudy (anonymous):

Can someone teach me how to multiply this: (2x + 5y)(3x^2 - 4xy + 2y^2)

OpenStudy (perl):

$$ \Large (2x + 5y)(3x^2 - 4xy + 2y^2) \\ \Large 2x \cdot 3x^2 + 2x \cdot (-4xy) + 2x \cdot 2y^2 + 5y \cdot 3x^2 +5y \cdot(- 4xy) + 5y \cdot 2y^2 \\ $$

OpenStudy (perl):

$$ \Large (2x + 5y)(3x^2 - 4xy + 2y^2) \\ \Large = 2x \cdot 3x^2 + 2x \cdot (-4xy) +2x \cdot 2y^2 \\ \Large + 5y \cdot 3x^2 +5y \cdot(- 4xy) + 5y \cdot 2y^2 \\ $$

OpenStudy (anonymous):

from there, would I combine like terms?